NONPARAMETRIC METHODS FOR CHECKING THE VALIDITY OF PRIOR ORDER INFORMATION

GUOHUA PAN

Department of Mathematics and Statistics, Oakland University, Rochester, MI 48309, U.S.A.

(Received August 17, 1998; revised June 7, 1999)

Abstract. A large number of statistical procedures have been proposed in the literature to explicitly utilize available information about the ordering of treatment effects at increasing treatment levels. These procedures are generally more efficient than those ignoring the order information. However, when the assumed order information is incorrect, order restricted procedures are inferior and, strictly speaking, invalid. Just as any statistical model needs to be validated by data, order information to be used in a statistical analysis should also be justified by data first. A common statistical format for checking the validity of order information is to test the null hypothesis of the ordering representing the order information. Parametric tests for ordered null hypotheses have been extensively studied in the literature. These tests are not suitable for data with nonnormal or unknown underlying distributions. The objective of this study is to develop a general distribution-free testing theory for ordered null hypotheses based on rank order statistics and score generating functions. Sufficient and necessary conditions for the consistency of the proposed general tests are rigorously established.

Key words and phrases: Distribution-free test, lack-of-fit, ordered null hypothesis, order restricted inferences, partial order.